the legs of a right triangle measure 5 inches and 7 inches


1
Expert's answer
2022-03-01T15:34:42-0500

Hypotenuse is "\sqrt{5^2+7^2}=\sqrt{74}." If "\u03b8" is the angle between the 7-inch leg and the hypotenuse then


"\cos \theta=\dfrac{7}{\sqrt{74}}\approx0.8137, \sin\theta=\dfrac{5}{\sqrt{74}}\approx 0.5812"